OpenStudy (anonymous):

Suppose that y varies inversely with x, and y = 4 when x = 7. What is an equation for the inverse variation

OpenStudy (anonymous):

if it varies inversely , that means that as y increases, then x decreases. What you have to if find the constant which is basically HOW MUCH x decreases by if y increases To do this, you construct the formula based on the given information that it is "inversely proportional" ---> y=1/x where y=k/x and k=constant so use the given values: 4=k/7, so k=28